OpenStudy (anonymous):

use the concept of rise over run to find the slope of the line. What is the slope of (0,1), (-3,7)

OpenStudy (akashdeepdeb):

Rise over run actually means how much you go/down up the y axis and then how much you go left/right on the x axis... AKA Change in y over change in x. Change in y = y2-y1 Change in x = x2-x1 from the points you are given okay? Understood? :)

OpenStudy (akashdeepdeb):

@J_Ballher44 ?

OpenStudy (anonymous):

=) thank you!!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."